# A linear independence criterion for certain infinite series with polynomial orders

### v1

@@ 1 Introduction and main results
In 1957, P. Erdős [7] proved that if $\ell\geq 1$ is an integer and $1<n_{1}<n_{2}<\cdots$ are integers with $\limsup_{k\to\infty}n_{k}/k^{\ell}=\infty$, then for any integer $b\geq 2$, the number $\sum_{k=1}^{\infty}1/b^{n_{k}}$ is transcendental or an algebraic number with degree at least $\ell+1$. For example, we see that if $f(x)$ is an integer-valued polynomial of degree $d\geq 2$ with $1<f(1)<f(2)<\cdots$, then for any integer $b\geq 2$, the number $\sum_{n=1}^{\infty}1/b^{f(n)}$ is transcendental or an algebraic number with degree at least $d$.
On the other hand, in 2019, V. Kumar [9] obtained linear independence result for certain infinite series with monomial orders. More precisely, he proved that if $k\geq 2$ and $1\leq a_{1}<a_{2}<\dots<a_{m}$ are integers with $k]{a_{i}/a_{j}}\notin\mathbb{Q}$ $(i\neq j)$, then for any integer $b\geq 2$, the numbers
@@
are linearly independent over $\mathbb{Q}$. Recently, S. Murakami and Y. Tachiya [11] gave a generalization of Kumar’s result by showing that, for any integer $b\geq 2$, the set of the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{1}{b^{in^{j}}}\quad(i=1,2,\dots,\ j=2,3,\dots)$
@@
In this paper, we study linear independence of the values of certain infinite series with polynomial orders, and as a corollary, get a generalization of linear independence result of Murkami and Tachiya [11]. Moreover, we give a linear independence criterion under specific conditions.
@@
D. Bertrand [1], and independently D. Duverney, Ke. Nishioka, Ku. Nishioka and I. Shiokawa [2] proved that for any algebraic number $\alpha$ $(0<|\alpha|<1)$ and any integer $m$ $(2\leq m\leq 4)$, the three values $\theta_{m}(\alpha)$, $\mathcal{D}\theta_{m}(\alpha)$ and $\mathcal{D}^{2}\theta_{m}(\alpha)$ are algebraically independent over $\mathbb{Q}$, where $\mathcal{D}\coloneqq z(d/dz)$ is a differential operator, by applying results of Yu. V. Nesterenko [12]. On the other hand, for any above numbers $\alpha$ and $m$, the four values $\theta_{m}(\alpha)$, $\mathcal{D}\theta_{m}(\alpha)$, $\mathcal{D}^{2}\theta_{m}(\alpha)$ and $\mathcal{D}^{3}\theta_{m}(\alpha)$ are algebraically dependent over $\mathbb{Q}$ since $\theta_{m}$ has an algebraic differential equation of the third order over $\mathbb{Q}$ (cf. [8]). In addition, Bertrand showed that for any algebraic number $\alpha$ $(0<|\alpha|<1)$, any two values among the three values $\theta_{2}(\alpha)$, $\theta_{3}(\alpha)$ and $\theta_{4}(\alpha)$ are algebraically independent over $\mathbb{Q}$. Conversely, these three values are algebraically dependent over $\mathbb{Q}$ by the Jacobi identity $\theta_{3}^{4}(z)=\theta_{2}^{4}(z)+\theta_{4}^{4}(z)$ (cf. [10]). By the way, C. Elsner and V. Kumar [6] recently derived linear independence results for functions and values related to $\theta_{3}$.
@@
This generalizes a result of Murakami and Tachiya [11, Corollary 1.3].
@@
Let $q$ be a Pisot or Salem number. The sequences $\{b_{j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell)$ are called linearly dependent over $\mathbb{Q}(q)$ if there exist algebraic numbers $k_{j}$ $(j=1,2,\dots,\ell)$ in $\mathbb{Q}(q)$ not all zero such that $\sum_{j=1}^{\ell}k_{j}\cdot b_{j}(n)=0$ for every integer $n\geq 0$. Otherwise, the sequences $\{b_{j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell)$ are called linearly independent over $\mathbb{Q}(q)$. In addition, we define an equivalence relation $\sim$ on the set of all polynomials with rational coefficients. For two polynomials $F(x)$ and $G(x)$ with rational coefficients, we write $F\sim G$ if there exist rational numbers $b>0$, $c$ and an integer $d$ such that
@@
Let $f_{j}(x)$ $(j=1,2,\dots,\ell)$ be polynomials with rational coefficients and $f_{j_{1}}\sim f_{j_{2}}$ for any integers $j_{1}$, $j_{2}$ $(1\leq j_{1}\leq j_{2}\leq\ell)$. Then, there exists a polynomial $g(x)$ such that there exist integers $B_{j}>0$, $C_{j}$ and $D_{j}$ $(j=1,2,\dots,\ell)$ satisfying
@@
Let $\mathcal{U}\coloneqq\bigcup_{i=1}^{m}\{(i,j)\in\mathbb{Z}_{\geq 1}^{2}\mid j=1,2,\dots,\ell_{i}\}$ and $f_{i,j}(x)$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients such that
$\begin{cases}f_{i_{1},j_{1}}\sim f_{i_{2},j_{2}}&\text{ if }\ i_{1}=i_{2},\\ f_{i_{1},j_{1}}\not\sim f_{i_{2},j_{2}}&\text{ if }\ i_{1}\neq i_{2}.\end{cases}$
@@
[…]
are linearly independent over $\mathbb{Q}(q)$.
In Theorem 1.2, if $\deg P_{i,j}(x)$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ are distinct, we immediately see that for each $i=1,2,\dots,m$, the sequences $\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly independent over $\mathbb{Q}(q)$.
@@
We give another example of Theorem 1.2 in Section 5.
The present paper is organized as follows. We will prove Theorem 1.1 in Section 3, and Theorem 1.2 in Section 4. Before proving Theorem 1.1, we give Lemma 2.1 in Section 2. Lemma 2.1 is the most important part to show Theorem 1.1. Also, we will prove Theorem 1.2 by using a result of Theorem 1.1. In Section 5, we give some examples of the set of polynomials which Theorems 1.1 and 1.2 are useful.
2 Lemma
In this section, we show the following lemma which plays an important role in the proof of Theorem 1.1. Lemma 2.1 is the following.
@@ 2 Lemma
for any integers $j\geq 2$, $C$, $D$ and any positive rational number $B$. Let $E$ be an arbitrary positive integer and let $G(x)$ be any real-valued function for $x\geq 1$ with $G(x)\to\infty$ $(x\to\infty)$ and $G(x)=o(x)$. Then, there exist infinitely many positive integers $m$ with $m\equiv A\pmod{E}$ such that
@@
$g(x)\coloneqq\max_{\begin{subarray}{c}y\\ 1\leq y\leq x\end{subarray}}G(y)\quad(x\geq 1).$
@@
$\mathcal{M}_{n}(j)\coloneqq\{m\in\mathcal{M}_{n}\mid\text{ there exists a positive integer $k_{m}$ such that }|f_{1}(m)-f_{j}(k_{m})|\leq g(n)\}.$
@@
#$\#\mathcal{M}_{n}(j_{0})>d\bigl{(}2g(n)+1\bigr{)}.$ (2.5)
@@
$|f_{1}(m)-f_{j_{m}}(k_{m})|\leq G(m)\leq\max_{\begin{subarray}{c}y\\ 1\leq y\leq n\end{subarray}}G(y)=g(n).$ (2.6)
@@
Then by (2.6), (2.7) and (2.8), we get $j_{m}\neq 1$ for any integer $m\in\mathcal{M}_{n}$, and hence we have $\ell\geq 2$. Thus by (2.4) and (2.6), we obtain
#$\#\bigcup_{j=2}^{\ell}\mathcal{M}_{n}(j)=\#\mathcal{M}_{n}\geq\frac{h(n)}{E_{0}}-1>d(\ell-1)\bigl{(}2g(n)+1\bigr{)}.$
Therefore, for every sufficiently large integer $n$, there exists an integer $j(n)$ $\bigl{(}2\leq j(n)\leq\ell\bigr{)}$ such that
#$\#\mathcal{M}_{n}\bigl{(}j(n)\bigr{)}>d\bigl{(}2g(n)+1\bigr{)}.$
Thus, we see that (2.5) holds.
@@
#$\#\mathcal{M}_{N}(j_{0})>d\bigl{(}2g(N)+1\bigr{)}.$ (2.9)
@@
If $d_{0}\geq d+1$, we get $f^{\prime}_{j_{0}}(\beta)/f^{\prime}_{1}(\alpha)\to\infty$ $(N\to\infty)$. This contradicts (2.16) since $w\geq 1$. Thus $d_{0}=d$, and hence $f^{\prime}_{j_{0}}(\beta)/f^{\prime}_{1}(\alpha)$ is bounded. Then by (2.16), we see that $w$ is a bounded positive integer.
@@
$E_{0}v=w\frac{\psi_{d}}{\lambda_{d}}\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{d-1}{d}}\left(1+O\Bigl{(}\frac{1}{N}\Bigr{)}\right)+O\left(\frac{1}{N}\right)+O\left(\frac{g(N)}{N^{d-1}}\right)=w\left(\frac{\psi_{d}}{\lambda_{d}}\right)^{\frac{1}{d}}+O\left(\frac{g(N)}{N}\right).$
@@
Step 5. Finally, we show that (2.2) contradicts (2.1). By (2.9), (2.20) and (2.21), there exists an integer $D_{0}$ with $|D_{0}|\leq g(N)$ such that the following equation
@@
$h(x)\coloneqq 1+\max_{\begin{subarray}{c}n\\ x\geq n\in\mathbb{Z}_{\geq 1}\end{subarray}}\left\{\sum_{j\in\mathcal{R}}\sum_{\sigma}\log_{q}(1+|a_{j}(n)^{\sigma}|)\right\}\quad(x\geq 1),$ (3.4)
@@
$h(N_{m,j}+n)\leq h\bigl{(}m\cdot H(m)^{\frac{1}{2}}+n\bigr{)}\leq\frac{m\cdot H(m)^{\frac{1}{2}}+n}{H\bigl{(}m\cdot H(m)^{\frac{1}{2}}+n\bigr{)}}\leq\frac{m\cdot H(m)^{\frac{1}{2}}+n}{H(m)}\leq m\cdot H(m)^{-\frac{1}{2}}+n.$
@@
Let $c_{1},c_{2},\dots$ be some positive constants independent of $m$. By (3.3), we get $|a_{1}(m)|>c_{1}$. Also by (3.4), for every $j\in\mathcal{R}$ and every embedding $\sigma\colon\mathbb{Q}(q)\to\mathbb{C}$, we obtain
@@
$1\leq|N_{\mathbb{Q}(q)/\mathbb{Q}}\ P_{m}(q)|$ $\leq c_{3}q^{-G(m)\left(1-H(m)^{-1/4}\right)}\left(c_{5}mq^{G(m)\cdot H(m)^{-1/4}}\right)^{[\mathbb{Q}(q):\mathbb{Q}]-1}$
@@
[…]
@@
is an integer-valued polynomial of degree $\geq 2$ with positive leading coefficients. Also, for each $i=1,2,\dots,m$, we obtain
@@
for any integers $10\leq r_{1},\ r_{2}\leq K_{i}-$ 1 $(r_{1}\neq r_{2})$, $A$, $C$, $D$ and any positive rational number $B$. Indeed, suppose to the contrary that there exist integers $i$ $(1\leq i\leq m)$, $r_{1}$, $r_{2}$ $(0\leq r_{1},\ r_{2}\leq K_{i}-1,$ $r_{1}\neq r_{2})$, $A$, $C$, $D$ and a positive rational number $B$ such that
@@
Then, we get $B=1$, $AK_{i}+r_{1}=CK_{i}+r_{2}$ and $D=0$. In particular, we get $r_{1}-r_{2}=K_{i}(C-A)$, and this contradicts $10\leq r_{1},\ r_{2}\leq K_{i}-$ 1 and $r_{1}\neq r_{2}$. Thus, we see that (4.4) holds. In addition, we have
@@
for every integers $10\leq r\leq K_{i}-$ 1 and $n\geq 0$. Therefore, the numbers (1.7) in Theorem 1.2 are linearly independent over $\mathbb{Q}(q)$ if and only if for each $i=1,2,\dots,m$, the sequences $\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly independent over $\mathbb{Q}(q)$. The proof of Theorem 1.2 is completed. ∎

### v2

@@ 1 Introduction and main results
In 1957, Erdős [7] proved that if $\ell\geq 1$ is an integer and $1<n_{1}<n_{2}<\cdots$ are integers with $\limsup_{k\to\infty}n_{k}/k^{\ell}=\infty$, then for any integer $b\geq 2$, the number $\sum_{k=1}^{\infty}1/b^{n_{k}}$ is transcendental or an algebraic number with degree at least $\ell+1$. For example, we see that if $f(x)$ is an integer-valued polynomial of degree $d\geq 2$ with $1<f(1)<f(2)<\cdots$, then for any integer $b\geq 2$, the number $\sum_{n=1}^{\infty}1/b^{f(n)}$ is transcendental or an algebraic number with degree at least $d$.
On the other hand, Kumar [10] obtained linear independence result for certain infinite series with monomial orders. More precisely, he proved that if $k\geq 2$ and $1\leq a_{1}<a_{2}<\dots<a_{m}$ are integers with $\sqrt[k]{a_{i}/a_{j}}\notin\mathbb{Q}$ $(i\neq j)$, then for any integer $b\geq 2$, the numbers
@@
are linearly independent over $\mathbb{Q}$. Recently, Murakami and Tachiya [12, 13] gave a generalization of Kumar’s result. In particular, they showed in [12, Corollary 1.3] that for any integer $b\geq 2$, the set of the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{1}{b^{in^{j}}}\quad(i=1,2,\dots,\ j=2,3,\dots)$ (1.1)
@@
In this paper, we study linear independence of certain infinite series with polynomial orders, and as a corollary, get a generalization of the linear independence result for the numbers (1.1). Moreover, we give a linear independence criterion under specific conditions.
@@
Let us observe the condition (i) in Theorem 1.1. Define $f(x)\coloneqq x^{2}$, $g(x)\coloneqq(x+1)^{2}$ and $h(x)\coloneqq(2x)^{2}$. Clearly, we have
$f(x+A)$ $\equiv g(x+A-1),$
$g(x+A)$ $\equiv f(x+A+1)$
for any integer $A$, and so the polynomials $f$ and $g$ do not satisfy the condition (i). On the other hand, for the polynomials $f$ and $h$, we have $f(x)\equiv h(x/2)$, while
$f(x+1)\not\equiv h(Bx+C)+D$
for any positive rational number $B$ and any integers $C$, $D$. Hence, the polynomials $f$ and $h$ satisfy the condition (i).
@@
Bertrand [1], and independently Duverney, Ke. Nishioka, Ku. Nishioka and
Shiokawa [2] proved that for any algebraic number $\alpha$ $(0\!<\!|\alpha|\!<\!1)$ and any integer $m$ $(2\leq m\leq 4)$,
the three values $\theta_{m}(\alpha)$, $\mathcal{D}\theta_{m}(\alpha)$ and $\mathcal{D}^{2}\theta_{m}(\alpha)$ are algebraically independent over $\mathbb{Q}$, where $\mathcal{D}\coloneqq z(d/dz)$ is a differential operator, by applying results of Nesterenko [14]. On the other hand, for any above numbers $\alpha$ and $m$, the four values $\theta_{m}(\alpha)$, $\mathcal{D}\theta_{m}(\alpha)$, $\mathcal{D}^{2}\theta_{m}(\alpha)$ and $\mathcal{D}^{3}\theta_{m}(\alpha)$ are algebraically dependent over $\mathbb{Q}$ since $\theta_{m}$ has an algebraic differential equation of the third order over $\mathbb{Q}$ (cf. [8]). In addition, Bertrand showed that for any algebraic number $\alpha$ $(0<|\alpha|<1)$, any two values among the three values $\theta_{2}(\alpha)$, $\theta_{3}(\alpha)$ and $\theta_{4}(\alpha)$ are algebraically independent over $\mathbb{Q}$. Conversely, these three values are algebraically dependent over $\mathbb{Q}$ by the Jacobi identity $\theta_{3}^{4}(z)=\theta_{2}^{4}(z)+\theta_{4}^{4}(z)$ (cf. [11]). By the way, Elsner and Kumar [6] recently derived linear independence results for functions and values related to $\theta_{3}$.
@@
This generalizes a result of Karmakar, Kumar and Thangadurai [9, Theorem 1.2].
@@
Let $q$ be a Pisot or Salem number. The sequences $\{b_{j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell)$ of complex numbers are called linearly dependent over $\mathbb{Q}(q)$ if there exist algebraic numbers $k_{j}$ $(j=1,2,\dots,\ell)$ in $\mathbb{Q}(q)$ not all zero such that $\sum_{j=1}^{\ell}k_{j}\cdot b_{j}(n)=0$ for every integer $n\geq 0$. Otherwise, the sequences $\{b_{j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell)$ are called linearly independent over $\mathbb{Q}(q)$. In addition, we define an equivalence relation $\sim$ on the set of all polynomials with rational coefficients. For two polynomials $F(x)$ and $G(x)$ with rational coefficients, we write $F\sim G$ if there exist rational numbers $b>0$, $c$ and an integer $d$ such that
@@
Let $f_{j}(x)$ $(j=1,2,\dots,\ell)$ be polynomials with rational coefficients such that $f_{j_{1}}\sim f_{j_{2}}$ for any integers $j_{1}$, $j_{2}$ $(1\leq j_{1}\leq j_{2}\leq\ell)$. Then, there exists a polynomial $g(x)$ such that there exist integers $B_{j}>0$, $C_{j}$ and $D_{j}$ $(j=1,2,\dots,\ell)$ satisfying
@@
Let $\ell_{1},\ell_{2},\dots,\ell_{m}$ be positive integers. Let
$\mathcal{U}\coloneqq\bigcup_{i=1}^{m}\{(i,j)\in\mathbb{Z}_{\geq 1}^{2}\mid j=1,2,\dots,\ell_{i}\}$
and $f_{i,j}(x)$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients such that
[…]
@@
$\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly independent over $\mathbb{Q}(q)$, where
$\tilde{p}_{i,j}(n)\coloneqq\begin{dcases*}P_{i,j}\left(\frac{n+C_{i,j}}{B_{i,j}}\right)&if $\ n\equiv-C_{i,j}\pmod{B_{i,j}}$,\\ 0&otherwise\end{dcases*}\qquad(j=1,2,\dots,\ell_{i}).$
In Theorem 1.2, if $\deg P_{i,j}(x)$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ are distinct, we immediately see that for each $i=1,2,\dots,m$, the sequences $\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly independent over $\mathbb{Q}(q)$. Hence, we have
@@
As an application of Theorem 1.2, we will give an example of linear dependence (see Example 5.2 in Section 5).
The present paper is organized as follows. We will prove Theorem 1.1 in Section 3, and Theorem 1.2 in Section 4. Before proving Theorem 1.1, we give Lemma 2.1 in Section 2. Lemma 2.1 is the most important part to show Theorem 1.1. Also, we will prove Theorem 1.2 by using a result of Theorem 1.1. In Section 5, we give some examples of the set of polynomials which Theorems 1.1 and 1.2 are useful. Throughout this paper, the term increasing function refers to a non-decreasing function.
2 A lemma
In this section, we show the following lemma which plays an important role in the proof of Theorem 1.1.
@@ 2 A lemma
for any integers $j\geq 2$, $C$, $D$ and any positive rational number $B$. Let $E$ be an arbitrary positive integer and let $G(x)$ be any real-valued function for $x\geq 1$ with $G(x)=o(x)$ and $G(x)\to\infty$ $(x\to\infty)$. Then, there exist infinitely many positive integers $m$ with $m\equiv A\pmod{E}$ such that
@@
$g(x)\coloneqq\max_{\begin{subarray}{c}y\in\mathbb{Z}\\ 1\leq y\leq x\end{subarray}}G(y)\quad(x\geq 1).$
@@
[…]
@@
$\mathcal{M}_{n}(j_{0})>d\bigl{(}2g(n)+1\bigr{)}.$ (2.5)
@@
$|f_{1}(m)-f_{j_{m}}(k_{m})|\leq G(m)\leq\max_{\begin{subarray}{c}y\in\mathbb{Z}\\ 1\leq y\leq n\end{subarray}}G(y)=g(n).$ (2.6)
@@
By (2.6), (2.7) and (2.8), we get $j_{m}\neq 1$ for any integer $m\in\mathcal{M}_{n}$, and hence we have $\ell\geq 2$. Then by (2.4) and (2.6), we obtain
$\bigcup_{j=2}^{\ell}\mathcal{M}_{n}(j)=\#\mathcal{M}_{n}\geq\frac{h(n)}{E_{0}}-1>d(\ell-1)\bigl{(}2g(n)+1\bigr{)}.$
Thus, for every sufficiently large integer $n$, there exists an integer $j(n)$ $\bigl{(}2\leq j(n)\leq\ell\bigr{)}$ such that $\mathcal{M}_{n}\bigl{(}j(n)\bigr{)}>d\bigl{(}2g(n)+1\bigr{)}$. Therefore, we see that (2.5) holds.
@@
$\mathcal{M}_{N}(j_{0})>d\bigl{(}2g(N)+1\bigr{)}.$ (2.9)
@@
If $d_{0}\geq d+1$, we get $f^{\prime}_{j_{0}}(\beta)/f^{\prime}_{1}(\alpha)\to\infty$ $(N\to\infty)$. This contradicts (2.16) since $w\geq 1$. Thus $d_{0}=d$, and hence $f^{\prime}_{j_{0}}(\beta)/f^{\prime}_{1}(\alpha)\asymp 1$. Then by (2.16), we see that $w$ is a bounded positive integer.
@@
$E_{0}v$ $=w\frac{\psi_{d}}{\lambda_{d}}\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{d-1}{d}}\left(1+O\Bigl{(}\frac{1}{N}\Bigr{)}\right)+O\left(\frac{1}{N}\right)+O\left(\frac{g(N)}{N^{d-1}}\right)$
$=w\left(\frac{\psi_{d}}{\lambda_{d}}\right)^{\frac{1}{d}}+O\left(\frac{g(N)}{N}\right).$
@@
Step 5. Finally, we show that (2.2) contradicts (2.1). By (2.9), (2.20) and (2.21), there exists an integer $D_{0}$ with $|D_{0}|\leq g(N)$ such that the equation
@@
$h(x)\coloneqq 1+\max_{\begin{subarray}{c}n\in\mathbb{Z}\\ 1\leq n\leq x\end{subarray}}\left\{\sum_{j\in\mathcal{R}}\sum_{\sigma}\log_{q}(1+|a_{j}(n)^{\sigma}|)\right\}\quad(x\geq 1),$ (3.4)
@@
$h(N_{m,j}+n)$ $\leq h\bigl{(}m\cdot H(m)^{\frac{1}{2}}+n\bigr{)}\leq\frac{m\cdot H(m)^{\frac{1}{2}}+n}{H\bigl{(}m\cdot H(m)^{\frac{1}{2}}+n\bigr{)}}\leq\frac{m\cdot H(m)^{\frac{1}{2}}+n}{H(m)}$
$\leq m\cdot H(m)^{-\frac{1}{2}}+n.$
@@
Let $c_{1},c_{2},\dots$ be some positive constants independent of $m$. By (3.3), we obtain $|a_{1}(m)|>c_{1}$. Also by (3.4), for every $j\in\mathcal{R}$ and every embedding $\sigma\colon\mathbb{Q}(q)\to\mathbb{C}$, we get
@@
$1\leq|N_{\mathbb{Q}(q)/\mathbb{Q}}(P_{m}(q))|$ $\leq c_{3}q^{-G(m)\left(1-H(m)^{-1/4}\right)}\left(c_{5}mq^{G(m)\cdot H(m)^{-1/4}}\right)^{[\mathbb{Q}(q):\mathbb{Q}]-1}$
@@
$\tilde{p}_{i,j}(K_{i}n+r)=\begin{dcases*}P_{i,j}\left(\frac{K_{i}}{B_{i,j}}n+\frac{r+C_{i,j}}{B_{i,j}}\right)&if $\ r\equiv-C_{i,j}\pmod{B_{i,j}}$,\\ 0&otherwise,\end{dcases*}$ (4.1)
@@
is an integer-valued polynomial of degree $\geq 2$ with a positive leading coefficient. Also, for each $i=1,2,\dots,m$, we obtain
@@
for any integers $0\leq r_{1},\ r_{2}\leq K_{i}-1$ $(r_{1}\neq r_{2})$, $A$, $C$, $D$ and any positive rational number $B$. Indeed, suppose to the contrary that there exist integers $i$ $(1\leq i\leq m)$, $r_{1}$, $r_{2}$ $(0\leq r_{1},\ r_{2}\leq K_{i}-1,$ $r_{1}\neq r_{2})$, $A$, $C$, $D$ and a positive rational number $B$ such that
@@
Then, we get $B=1$, $AK_{i}+r_{1}=CK_{i}+r_{2}$ and $D=0$. In particular, we get $r_{1}-r_{2}=K_{i}(C-A)$, and this contradicts $0\leq r_{1},\ r_{2}\leq K_{i}-1$ and $r_{1}\neq r_{2}$. Thus, we see that (4.4) holds. In addition, we have
@@
for every integers $0\leq r\leq K_{i}-1$ and $n\geq 0$. Therefore, if the numbers (1.8) in Theorem 1.2 are linearly dependent over $\mathbb{Q}(q)$, then, for each $i=1,2,\dots,m$, the sequences $\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly dependent over $\mathbb{Q}(q)$. Conversely, assuming for each $i=1,2,\dots,m$ that $\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly dependent over $\mathbb{Q}(q)$, then we can show by (4.3) that the numbers (1.8) in Theorem 1.2 are linearly dependent over $\mathbb{Q}(q)$. The proof of Theorem 1.2 is completed. ∎